# SMARTER: A Data‑efficient Framework to Improve Toxicity Detection with Explanation via Self-augmenting Large Language Models

### v1

@@ Abstract
WARNING: This paper contains examples of offensive materials. Toxic content has become pervasive on social media platforms. We introduce SMARTER, a data-efficient 2-stage framework for explainable content moderation using Large Language Models (LLMs). In Stage 1, we leverage LLMs’ own outputs to generate synthetic explanations for both correct and incorrect labels, enabling alignment via preference optimization with minimal human supervision. In Stage 2, we refine explanation quality through cross-model training, allowing weaker models to align stylistically and semantically with stronger ones. Experiments on 3 benchmark tasks (HateXplain, Latent Hate, and Implicit Hate) demonstrate that SMARTER enables LLMs to achieve up to 13.5% macro-F1 improvement over standard few-shot baselines with a fraction of the full training data. Our framework offers a scalable strategy for low-resource settings by harnessing LLMs’ self-improving capabilities for both classification and explanation.
@@ 1 Introduction
While capable of achieving impressive classification performance, these increasingly sophisticated models require ample resources to train, especially when the target label space cover nuanced concepts Fortuna et al. (2021); Sap et al. (2021); Zhao et al. (2021). Developing a robust dataset for training typically demands significant effort in curation and human annotation, often involving iterative updates to adapt the model to evolving linguistic trends Toraman et al. (2022); Bespalov et al. (2023); Rad et al. (2025). In addition, this pipeline is limited to decision only. Failure to provide users with human-understandable signals renders the adjudication process opaque, potentially inviting further issues on transparency and trustworthiness.
@@
    We introduce SMARTER [^1] Our repository can be found at https://github.com/hnghiem-nlp/smarter_public (Self-augMentAtion Regimen Towards Efficient Content ModeRation), a 2-stage framework where (1) each LLM self-augments on few-shot data using preference optimization with synthetic explanations, (2) cross-model refinement to further improvement by training one LLM with responses generated by another on held-out sets.
@@
    Experiments are conducted on 3 benchmark datasets (HateXplain, Latent Hate, Implicit Hate) using open-source LLMs (Llama-3.1-8B-Instruct and COT-T5-XL), in few-shot settings and explainable outputs.
Our results show that leveraging multiple LLMs enables high-quality, explainable content moderation even with limited training data. SMARTER thus provides a scalable and practical approach for low-resource NLP practitioners to achieve both strong performance and explanability.
@@ 2 Related Works
Given the diverse definitions of toxicity, prior work has explored cross-domain robustness via data augmentation and architectural innovations Caselli et al. (2021); Sarkar et al. (2021); Pan et al. (2020); Toraman et al. (2022). Antypas and Camacho-Collados (2023) show that training on combined datasets improves generalization, while Nghiem et al. (2024) demonstrate that incorporating label definitions supports low-shot adaptation in BERT-based models.
Recent work has begun leveraging LLMs for content moderation, given their strong zero- and few-shot performance on toxic classification tasks Masud et al. (2024); Jahan et al. (2024); Roy et al. (2023); Kumarage et al. (2024). Beyond accuracy, LLMs can generate textual explanations, offering interpretable signals for moderators and users alike Di Bonaventura et al. (2024); Yang et al. (2023). Leveraging these insights, we propose a framework that trains open-source LLMs to jointly perform classification and explanation across toxic domains, enhancing transparency in moderation workflows.
@@
Despite their impressive capabilities, LLMs still have their limitations: they may generate unhelpful responses to user queries Liu et al. (2023); Shen et al. (2023), or sycophantic outputs that pander to the user’s perceived preference when prompted Papadatos and Freedman (2024); Ranaldi and Pucci (2023) . Various techniques have been introduced to align LLMs towards preferred generative patterns: from finetuning approaches Dai et al. (2023); Sun et al. (2023); Wang et al. (2023) to interactive mechanisms (prompting) Furniturewala et al. (2024); Guo et al. (2022). A promising line of research showcases LLMs’ self-improving potential via their own generated synthetic data Tang et al. (2023); Li et al. (2023). Combining these approaches, we exploit LLMs’ sycophantic behavior to augment limited training data resource and improve downstream performance via alignment techniques.
@@ 3 Overview of Framework
As shown in Algorithm 1, our framework is organized into two stages. First, we focus on optimizing the LLMs’ classification performance through alignment tuning using self-augmenting synthetic data generation. In the second stage, we refine the quality of the generated explanations by cross-model refinement, leveraging the synergy between different LLMs. Section 4 and 5 present the experimental setup and findings for Stage 1. Section 6 focuses on cross-model improvement in Stage 2.
@@ 4.1 Data
Detailed definition of the terminologies associated with the labels are also included, allowing us to integrate these contexts into the explanation in downstream applications in a fashion similar to Yang et al. (2023) and Nghiem and Daumé (2024). We hereby refer to the 3 classification tasks based on their distinct label spaces of as HateXplain, Latent Hate, Implicit Hate respectively. Table 1 summarizes the main components of these datasets and the split for training, validation and testing. Appendix B describes these pre-processing steps.
Dataset Total Size Split Ratio K val K test Platform Label Space
HateXplain 20,148 60:20:20 50 400 G, T Normal, Offensive, Hate
@@ 4.2 Models
    COT-T5-XL: An encoder-decoder transformer-based Vaswani (2017) LLM introduced by Kim et al. (2023). A variant of the base Flan-T5-XL Longpre et al. (2023) architecture, this model consists of 3 billion parameters and is further trained on the COT corpus, a collection of 1.8 million samples with chain-of-thought style explanations, which enhances its capabilities on a variety of reasoning tasks Kim et al. (2023).
@@
For brevity, we refer to the models as T5 and Llama respectively for the remaining of this paper.
@@ 4.3 Classification via Prompting
As demonstrated by Wei et al. (2022), prompting an LLM of sufficient size to provide justification before the decision enhances its downstream performance. We apply this Chain-of-Thought (COT) paradigm in our classification tasks to obtain the explanation from the LLMs necessary for content moderation. Inspired by Nghiem and Daumé (2024) and Yang et al. (2023), we use the prompt shown in Figure 1, which provides the target post, the set of labels corresponding to the task and their associated definitions, and the instruction for the LLM to provide its explanation and predicted category using this context.
### Instruction:
Provide your message in the following format:
EXPLANATION: [text]
LABEL: [text]
### Definitions:
Hate: Language that targets [...]
Abusive: Language that is [...]
Normal: Does not contain [...]
### Post:
{post}
### Response: [model output]
@@ 4.4 Auxiliary Pretraining
Leveraging existing resources, we train both T5 and Llama models on HateCOT, a dataset introduced by Nghiem and Daumé (2024), which consists of toxic posts, their labels, definitions and synthetic explanations designed to boost of-the-shelf LLMs’ performance in hate speech detection. Specifically, the models are supervised finetuned on the next token prediction tasks using inputs constructed by combining HateCOT’s data with our prompt template. Finetuning is performed by training LoRA adapters Hu et al. (2021), a parameter-efficient approach implemented by the HuggingFace library (see Appendix E), which serve as the basis for downstream experiments.
@@ 5.1 K-shot In-domain Finetuning
We experiment with different configurations of limited in-domain data. For each task, we select uniformly at random K-shot $\in\{16,32,64,128,256\}$ samples from the training set . We utilize the seed explanations by Nghiem and Daumé (2024), which was generated by OpenAI’s GPT-3.5-Turbo. Further finetuning the existing HateCOT-pretrained T5 and Llama models happens in a fashion similar to the auxiliary pretraining phase above.
@@
    Direct Preference Optimization (DPO): An offline Reinforcement Learning (RL) technique that optimizes the policy via an implicit reward model using preference data Rafailov et al. (2024). DPO requires human preference data where a pair of chosen vs. rejected responses is presented for each prompt.
@@
Appendix E.2 offers more technical details on these techniques. Our self-augmenting methodology naturally procures training datasets in the required format for KTO by simply designating the explanation conditioned on the gold label as positive, and otherwise negative. For DPO, we match the explanations conditioned on the correct-incorrect pair sequentially to satisfy the pairwise chosen-rejected format. As an example, 100 posts of HateXplain (whose label space is of size 3), should produce 1003=300100*3=300 training samples for KTO, and 100(31)=200100*(3-1)=200 for DPO.
@@ 5.1.1 Experimental Pipeline
The experiments follow a two-stage process. First, we finetune T5 and Llama under extreme data scarcity by setting $K\in\{16,32,64,128\}$, followed by self-augmentation via DPO to compare classification performance. Then, we experiment with significantly more training data with $K=256$ and evaluate KTO alignment against DPO  [^2] Implementation for KTO trainer for T5 is not available at the time of writing, so we are only able to perform this variant for Llama. For comparison, we also finetune both LLMs with the entire available training data to obtain classification lable without explanation, denoted Full models. In addition, we also finetune ModernBERT (large) Warner et al. (2024), a decoder-only model on the full training data and obtain the classification outputs using the same test sets.
In addition, we examine 2 partial sampling strategies to circumvent the proportionate growth of self-augmenting synthetic data to label space when K is large . In the first, we construct augmented data by selecting 128 and 192 shots of samples uniformly at random from the Baseline’s 256-shot pool. In the second, instead of $K$-shot sampling, we select $K^{\prime}*|S|*(|S|-1)$ samples from all possible post-label pairs of the 256-shot Baseline pool (incorrect labels included) to create augmented data, where $K^{\prime}\in\{128,192\}$ and $|S|$ is the label space’ size. While both approaches should yield identical size of training data for the same $K^{\prime}$, the first method, denoted DPO-K, preserves post-label parity whereas the second method, denoted DPO-N, yields higher diversity of post content.
Experiments are implemented with the HuggingFace library. We sample $K=50$ shot on the held out Validation extraction of each dataset to tune hyperparamters, and evaluate on samples from the Test set as shown in Table 1. Technical details are included in Appendix E.
@@ 5.2.1 Results for $K<=128$
Notable is DPO self-augmentation’s ability to improve over Baselines at even $K\leq 64$, showcasing our method’s advantage in low-resource settings. Llama Baseline models appear to consistently outperform T5 across all tasks. However, with DPO self-augmentation, T5 is able to narrow the margin on Latent Hate and Implicit Hate. In contrast, Llama still dominates on HateXplain. These discrepancies suggest an interaction between dataset and model choice, and practitioners may benefit from contrasting different options with our self-augmenting techniques for best results.
@@ 5.2.2 Results for $K=256$
In Figure 3, we observe the macro F1 scores for models $K=256$ on the subfigures on the right. DPO-augmentation continues to enhance Baseline models which already benefits from more training data relative to lower $K$ values, with improvement in scores reaching up to 13.2% for LLama and 14% for T5. In contrast, KTO-augmentation appears ineffectual on HateXplain and even hinders performance on Latent Hate and Implicit Hate, showing notable reduction from Baseline scores, respectively. With this, we hypothesize that DPO’s contrastive nature reinforces signals that allow the the LLM to learn from its own mistakes, thereby sharpening its reasoning ability for this type of discriminatory task. On the other hand, KTO’s reliance on implicit signals from listwise data is not sufficient to improve discernment of classes, particular when the semantic distinction between labels are nuanced as in the case of Latent Hate and Implicit Hate.
@@
Dataset Model F1 Full F1% Data%
HateXplain Llama 0.72 89% 6%
T5 0.72 86%
ModernBERT 0.70 100% 100%
Latent Hate Llama 0.62 111% 7%
T5 0.61 107%
ModernBERT 0.61 100% 100%
Implicit Hate Llama 0.67 90% 57%
T5 0.64 92%
ModernBERT 0.64 100% 100%
Table 2: Comparison of performance and data usage of DPO-augmented vs. Full models at $K=256$. F1 Full denotes the macro F1 scores of Full models that use the entire available training data. F1% is the percentage of the F1 score attained by the corresponding DPO-augmented model over F1 Full. Data% is the percentage of size of the training data using $K=256$ shot over the full training samples. ModernBERT’s scores are presented for references.
Remarks
Our empirical results demonstrates that our DPO augmentation enables LLMs to achieve considerable performance with low resource. From Figure 3, DPO-augmented models attain high macro F1 scores at $K=64$, a very small amount of data. Table 2 further showcases that DPO-augmented models can achieve at least 86% of F1 performance (HateXplain) or even exceed that of their Full counterpart (Latent Hate) while using only fraction of the available training data at $K=256$. For practitioners, these insights suggest a resource-conservative strategy: start with low K-shot samples (64) to obtain benchmark results, then iteratively add more training samples until performance plateaus or resource is exhausted.
@@ 6.1 Human Evaluation
6.2 Cross-model Stylistic Refinement
@@
For each test domain, we additionally train a style classifier based on the BERT architecture using explanations generated individually by T5 and Llama of the SFT+DPO variant at $K=128$ Devlin et al. (2019). A classification head is added to BERT, which leverages the [CLS] token from each explanation’s embedding to predict whether it was produced by T5 or LLaMA. Technical details are reported in Appendix E.2.
6.2.1 Cross-model Refinement Results

### v2

@@ Abstract
WARNING: This paper contains examples of offensive materials. To address the proliferation of toxic content on social media, we introduce SMARTER, a data-efficient 2-stage framework for explainable content moderation using Large Language Models (LLMs). In Stage 1, we leverage LLMs’ own outputs to generate synthetic explanations for both correct and incorrect labels, enabling alignment via preference optimization with minimal human supervision. In Stage 2, we refine explanation quality through cross-model training, allowing weaker models to align stylistically and semantically with stronger ones. Experiments on 3 benchmark tasks (HateXplain, Latent Hate, and Implicit Hate) demonstrate that SMARTER enables LLMs to achieve up to 13.5% macro-F1 improvement over standard few-shot baselines with a fraction of the full training data. Our framework offers a scalable strategy for low-resource settings by harnessing LLMs’ self-improving capabilities for both classification and explanation.
@@ 1 Introduction
While capable of achieving impressive classification performance, these increasingly sophisticated models require ample resources to train, especially when the target label space covers nuanced concepts Fortuna et al. (2021); Sap et al. (2021); Zhao et al. (2021). Developing a robust dataset for training typically demands significant effort in curation and human annotation, often involving iterative updates to adapt the model to evolving linguistic trends Toraman et al. (2022); Bespalov et al. (2023); Rad et al. (2025). In addition, this pipeline is limited to decision-only pipeline. Failure to provide users with human-understandable signals renders the adjudication process opaque, inviting further issues on transparency and trustworthiness.
@@
    We introduce SMARTER [^1] Our repository can be found at REDACTED (Self-augMentAtion Regimen Towards Efficient Content ModeRation), a 2-stage framework where (1) each LLM self-augments on few-shot data using preference optimization with synthetic explanations, (2) cross-model refinement to further improvement by training one LLM with responses generated by another on held-out sets.
@@
    Experiments conducted on 3 benchmark datasets (HateXplain, Latent Hate, Implicit Hate) with Llama-3.1-8B-Instruct and COT-T5-XL demonstrate up to 13.5% increase in macro-F1 scores in few-shot settings with high-quality explanations, outperforming a range of commercial models.
SMARTER thus provides a scalable and practical approach for low-resource NLP practitioners to achieve strong performance and explainability.
@@ 2 Related Works
Existing research Caselli et al. (2021); Sarkar et al. (2021); Toraman et al. (2022) has proposed strategies to enhance cross-domain robustness in toxicity detection, such as combining multiple datasets to improve generalization Antypas and Camacho-Collados (2023) and incorporating label definitions to support few-shot adaptation Nghiem et al. (2024). Building on these advances, recent studies leverage LLMs for toxicity detection, demonstrating strong zero- and few-shot performance Masud et al. (2024); Jahan et al. (2024); Roy et al. (2023); Kumarage et al. (2024). LLMs can also generate textual explanations, enhancing interpretability for moderators and users Di Bonaventura et al. (2024); Yang et al. (2023). We extend this line of work by training open-source LLMs to jointly perform classification and explanation to promote transparency in moderation workflows.
@@
In addition to extensive pretraining, modern LLMs also undergo post-training that both unlocks additional capabilities Longpre et al. (2023) and aligns the outputs towards human preferences Ouyang et al. (2022). Recent Reinforcement Learning (RL)-based techniques like GRPO Shao et al. (2024); Zweiger et al. (2025) enable promising LLM self-alignment at expensive computational cost. Another body of work showcases LLMs to self-improve via their own synthetic data Tang et al. (2023); Li et al. (2023), such as Self-Refine Madaan et al. (2023) and Self-Instruct Wang et al. (2022). Building upon these ideas, our framework presents a simple yet principled approach that allows LLMs to iteratively improve with minimal training cost while still amenable to human supervision.
@@ 3 Overview of Framework
As shown in Algorithm 1, our framework consists of two stages. First, we focus on optimizing the LLMs’ classification performance through alignment tuning using self-augmenting synthetic data generation. In the second stage, we refine the quality of the generated explanations by cross-model refinement, leveraging the synergy between different LLMs. Section 4 and 5 present the experimental setup and findings for Stage 1. Section 6 focuses on cross-model improvement in Stage 2. Section 7 analyzes the consistency of explanations with respect to the definitions and labels.
@@ 4.1 Data
Definitions of the associated labels are also included, allowing us to integrate these contexts into the explanation in downstream applications in a fashion similar to Yang et al. (2023) and Nghiem and Daumé (2024). We refer to the three classification tasks by their label spaces as HateXplain, Latent Hate, and Implicit Hate. Table 4 summarizes their components and data splits, with preprocessing details in Appendix B.
@@ 4.2 Models
    COT-T5-XL: A variant of the base encoder-decoder Flan-T5-XL Longpre et al. (2023) architecture, this model consists of 3 billion parameters and is further trained on the COT corpus, a collection of 1.8 million samples with chain-of-thought style explanations, which enhances its capabilities on a variety of reasoning tasks Kim et al. (2023).
@@
For brevity, we refer to the models as T5 and Llama respectively for the remainder of this paper.
@@ 4.3 Classification via Prompting
We apply this Chain-of-Thought (COT) paradigm Wei et al. (2022) in our classification tasks to obtain the explanation from the LLMs necessary for content moderation. Inspired by Nghiem and Daumé (2024) and Yang et al. (2023), we use the prompt in Figure 6, which incorporates the target post, the set of labels corresponding to the task and their associated definitions as context for the LLM to provide its explanation and predicted category.
@@ 4.4 Auxiliary Pretraining
Nghiem and Daumé (2024) found that supervised finetuning (SFT) off-the-shelf LLMs on HateCOT – a general compilation of toxic posts, labels, definitions and synthetic explanations – can boost their performance in downstream hate speech detection. To maximize expected gains with minimal in-domain data, we adopt this strategy and perform SFT on the 2 chosen LLMs using HateCOT to prime them for our classification tasks. Finetuning is performed by training LoRA adapters Hu et al. (2021), a parameter-efficient approach implemented by the HuggingFace library (see Appendix F) as the basis for downstream experiments.
@@ 5.1 K-shot In-domain Finetuning
We experiment with different configurations of limited in-domain data. For each task, we select uniformly at random K-shot $\in\{16,32,64,128,256\}$ samples from the training set . We utilize the seed explanations by Nghiem and Daumé (2024) to further SFT the existing HateCOT-pretrained T5 and Llama models in a fashion similar to the auxiliary pretraining phase above.
@@
    Direct Preference Optimization (DPO): An offline RL-based technique that optimizes the policy via an implicit reward model using preference data Rafailov et al. (2024). DPO requires human preference data where a pair of chosen vs. rejected responses is presented for each prompt.
@@
Appendix F.3 offers more technical details on these techniques. Our self-augmenting methodology naturally procures training datasets in the required format for KTO by simply designating the explanation conditioned on the gold label as positive, and otherwise negative. For DPO, we match the explanations conditioned on the correct-incorrect pair sequentially to satisfy the pairwise chosen-rejected format. As an example, 100 posts of HateXplain (whose label space is of size 3), should produce 1003=300100*3=300 training samples for KTO, and 100(31)=200100*(3-1)=200 for DPO.
@@ 5.1.1 Experimental Pipeline
The experiments follow a two-stage process. First, we finetune T5 and Llama under extreme data scarcity by setting $K\in\{16,32,64,128\}$, followed by self-augmentation via DPO to compare classification performance. Then, we experiment with significantly more training data with $K=256$ and evaluate KTO alignment against DPO  [^2] Implementation for KTO trainer for T5 is not available at the time of writing, so we are only able to perform this variant for Llama. For comparison, we also finetune both LLMs with the entire available training data to obtain classification label without explanation, denoted Full models. In addition, we also finetune ModernBERT (large) Warner et al. (2024), a decoder-only model on the full training data and obtain the classification outputs using the same test sets.
In addition, we examine 2 partial sampling strategies to circumvent the proportionate growth of self-augmenting synthetic data to label space when K is large . In the first, we construct augmented data by selecting 128 and 192 shots of samples uniformly at random from the Baseline’s 256-shot pool. In the second, instead of $K$-shot sampling, we select $K^{\prime}*|S|*(|S|-1)$ samples from all possible post-label pairs of the 256-shot Baseline pool (incorrect labels included) to create augmented data, where $K^{\prime}\in\{128,192\}$ and $|S|$ is the label space’s size. While both approaches should yield identical size of training data for the same $K^{\prime}$, the first method, denoted DPO-K, preserves post-label parity whereas the second method, denoted DPO-N, yields higher diversity of post content.
Experiments are implemented with the HuggingFace library. We sample $K=50$ shot on the held out Validation extraction of each dataset to tune hyperparamters, and evaluate on samples from the Test set as shown in Table 4. Technical details are included in Appendix F.
@@ 5.2.1 Results for $K<=128$
Notably, DPO self-augmentation improves performance even at $K\leq 64$, showcasing our method’s advantage in low-resource settings. Llama Baseline models appear to consistently outperform T5 across all tasks. However, with DPO self-augmentation, T5 is able to narrow the margin on Latent Hate and Implicit Hate. In contrast, Llama still dominates on HateXplain. These discrepancies suggest an interaction between dataset and model choice, and practitioners may benefit from contrasting different options with our self-augmenting techniques for best results.
@@ 5.2.2 Results for $K=256$
In Figure 1, we observe the macro F1 scores for models $K=256$ on the subfigures on the right. DPO-augmentation continues to enhance Baseline models which already benefits from more training data relative to lower $K$ values, with improvement in scores reaching up to 13.2% for LLama and 14% for T5. In contrast, KTO-augmentation appears ineffectual on HateXplain and even hinders performance on Latent Hate and Implicit Hate, showing notable reduction from Baseline scores, respectively. With this, we hypothesize that DPO’s contrastive nature reinforces signals that allow the LLM to learn from its own mistakes, thereby sharpening its reasoning ability for this type of discriminatory task. On the other hand, KTO’s reliance on implicit signals from listwise data is not sufficient to improve discernment of classes when the semantic distinction between labels are nuanced for Latent Hate and Implicit Hate.
@@
Dataset Model F1 F1% Data%
HateXplain Llama_DPO 0.64 89% 6%
Llama_Full 0.72 100% 100%
T5_DPO 0.62 86% 6%
T5_Full 0.72 100% 100%
ModernBERT 0.70 97% 100%
GPT-4o-mini 0.50 69%
GPT-4.1 0.55 76%
GPT-5-chat 0.56 78%
Qwen-32B 0.54 75%
Latent Hate Llama_DPO 0.69 100% 7%
Llama_Full 0.62 90% 100%
T5_DPO 0.65 94% 7%
T5_Full 0.61 88% 100%
ModernBERT 0.61 88% 100%
GPT-4o-mini 0.54 78%
GPT-4.1 0.60 87%
GPT-5-chat 0.51 74%
Qwen-32B 0.47 68%
Implicit Hate Llama_DPO 0.60 90% 57%
Llama_Full 0.67 100% 100%
T5_DPO 0.59 88% 57%
T5_Full 0.64 96% 100%
ModernBERT 0.64 96% 100%
GPT-4o-mini 0.42 63%
GPT-4.1 0.57 85%
GPT-5-chat 0.58 87%
Qwen-32B 0.49 73%
Table 1: Comparison of performance and data usage across models (including additional GPT baselines). We report Macro F1 (F1), relative score (F1%), and training data fraction (Data%). Models with the DPO suffix are DPO-augmented with K=256K{=}256 shots, while Full uses the entire training set. Within each dataset, F1 is bolded for the best model, and F1% is normalized to that best model (100%). Data% is measured relative to the corresponding Full baseline (100%). ModernBERT is shown for reference.
Comparison against Alternative Baselines
We further evaluate zero-shot classification on three commercial OpenAI LLMs and Qwen-32B, alongside off-the-shelf T5 and Llama variants, using the prompt in Figure 5. As shown in Table 1 and Table 5, our DPO-augmented models at $K=256$ consistently outperform both their baselines and larger LLMs, demonstrating SMARTER’s accessibility advantage—achieving controllable, high-quality performance with lightweight, low-cost models that do not rely on proprietary platforms.
Following the same setup as ModernBERT, we also SFT-finetune base Llama and T5 for classification without explanations. At $K=256$, our DPO-augmented variants retain at least 86% of the Full model’s F1 on HateXplain and surpass it on Latent Hate, underscoring the efficiency and scalability of our low-resource alignment framework.
@@ 6.1 Human Evaluation
6.2 Cross-model Refinement
6.2.1 Cross-model Training Methodology
@@
6.2.2 Cross-model Classification Results